OpenStudy (anonymous):

Help please! Reduce to lowest terms: ((x^(2)+4x-5)/(x^(2)-3x-28)÷(x^(2)-25)/(x^(2)-49))*(x-5)/(x-1)

OpenStudy (campbell_st):

best advice is to factorise the fractions....

OpenStudy (anonymous):

File attached

OpenStudy (campbell_st):

here is the 1st fraction \[\frac{x^2 + 4x - 5}{x^2 - 3x - 28} = \frac{(x + 5)(x -1)}{(x - 7)(x + 4)}\] you need to repeat the process for the 2nd fraction. the once you do that, its order of operation look the cancel common factors and din't for get... dividing by a fraction, flip and multiply